OpenStudy (anonymous):

The youth group is going on a trip to the state fair. The trip costs $63. Included in that price is $13 for a concert ticket and the cost of 2 passes, one for the rides and one for the game booths. Each of the passes costs the same price. Write an equation representing the cost of the trip, and determine the price of one pass. Solve your equation by showing your work and steps.

OpenStudy (anonymous):

2x+13=63

OpenStudy (anonymous):

wait @starlord6200

OpenStudy (anonymous):

How did you come up with 2x+13+63?

OpenStudy (anonymous):

I said equals 63

OpenStudy (anonymous):

Yes 2x+13=63

OpenStudy (anonymous):

So that's the answer

OpenStudy (anonymous):

U still there

OpenStudy (anonymous):

no its the other way around.

OpenStudy (anonymous):

No it's that

OpenStudy (ytrewqmiswi):

ฏ๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎๎ํํํํํํํํํํํํํํํํํํํํํ¯\_Ỏ̷͖͈̞̩͎̻̫̫̜͉̠̫͕̭̭̫̫̹̗̹͈̼̠̖͍͚̥͈̮̼͕̠̤̯̻̥̬̗̼̳̤̳̬̪̹͚̞̼̠͕̼̠̦͚̫͔̯̹͉͉̘͎͕̼̣̝͙̱̟̹̩̟̳̦̭͉̮̖̭̣̣̞̙̗̜̺̭̻̥͚͙̝̦̲̱͉͖͉̰̦͎̫̣̼͎͍̠̮͓̹̹͉̤̰̗̙͕͇͔̱͕̭͈̳̗̭͔̘̖̺̮̜̠͖̘͓̳͕̟̠̱̫̤͓͔̘̰̲͙͍͇̙͎̣̼̗̖͙̯͉̠̟͈͍͕̪͓̝̩̦̖̹̼̠̘̮͚̟͉̺̜͍͓̯̳̱̻͕̣̳͉̻̭̭̱͍̪̩̭̺͕̺̼̥̪͖̦̟͎̻̰_Ỏ̷͖͈̞̩͎̻̫̫̜͉̠̫͕̭̭̫̫̹̗̹͈̼̠̖͍͚̥͈̮̼͕̠̤̯̻̥̬̗̼̳̤̳̬̪̹͚̞̼̠͕̼̠̦͚̫͔̯̹͉͉̘͎͕̼̣̝͙̱̟̹̩̟̳̦̭͉̮̖̭̣̣̞̙̗̜̺̭̻̥͚͙̝̦̲̱͉͖͉̰̦͎̫̣̼͎͍̠̮͓̹̹͉̤̰̗̙͕͇͔̱͕̭͈̳̗̭͔̘̖̺̮̜̠͖̘͓̳͕̟̠̱̫̤͓͔̘̰̲͙͍͇̙͎̣̼̗̖͙̯͉̠̟͈͍͕̪͓̝̩̦̖̹̼_/¯ค้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้้

OpenStudy (anonymous):

okay so the trip costs $63 all together, and this includes the $13 dollar concert ticket and two passes. we don't know how much these two passes are, so we will write it as 'p' (1) 13+2p=63 <-- this is the equation representing the cost of the trip (2) 2p=50 <-- minus 13 from both sides (3) p=25 <-- divide both dies by 2 so one pass is $25 This What I got ?

OpenStudy (anonymous):

@starlord6200

OpenStudy (anonymous):

Yea so either way it's the same thing but yes each ticket is 25